OpenStudy (anonymous):

How do I solve please help Two slices of bacon contain 85 calories. How many calories are there in 12 slices of bacon?

OpenStudy (campbell_st):

use multiplication 2 x ? = 12 whats the ?

OpenStudy (campbell_st):

so if 2 rashers = 85 calories 2 x ? rashers = 85 x ? calories for the

hartnn (hartnn):

2 slice= 85 cal 1slice = 85/2 cal 12 slices = (85/2) * 12 = ??